OpenStudy (anonymous):

divide simplify w^2+5w/w^2+2w-15 divide by w/w+3

jhonyy9 (jhonyy9):

w(w+5)/(w+5)(w-3)/w/w+3= w/w-3/w/w+3)=(w/w-3)*(w+3/w)=(w+3)/(w-3)